40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5870     Entry time: Fri Nov 11 00:58:19 2011
Author: Den 
Type: Update 
Category: elog 
Subject: restarted elog 

Elog suspended 2 times for 1 hour. Too high frequency.

Restarted.

ELOG V3.1.3-